OpenStudy (anonymous):

what is the slope of the line that passes through the pair of points (-5/3 -1) and (-2 9/2) choices : 2/33 -2/33 33/2 -33/2

OpenStudy (anonymous):

Here's the formula for the slope between two points (a,b) and (c,d): \[\textrm{slope} = \frac{d-b}{c-a}\] So, plug the numbers in: \[\textrm{slope} = \frac{d-b}{c-a} = \frac{\frac{9}{2}-(-1)}{(-2)-(-\frac{5}{3})} = \frac{\frac{11}{2}}{-\frac{1}{3}} = -\frac{33]{2}\] -33/2, or Choice D. Tell me if that all make sense.

OpenStudy (anonymous):

Uh oh, bad equation, lemme fix that.

OpenStudy (anonymous):

\[\textrm{slope} = \frac{d-b}{c-a} = \frac{\frac{9}{2}-(-1)}{(-2)-(-\frac{5}{3})} = \frac{\frac{11}{2}}{-\frac{1}{3}} = -\frac{33}{2}\]

OpenStudy (anonymous):

There, much better.